OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Convert the following equation from slope intercept form to standard form. In other words, if the equation is rewritten to look like Ax+By=C, what are the values of A, B, and C? Assume A is positive. y=−x−1 THANK YOU ! :)

OpenStudy (anonymous):

CAN YOU ALSO EXPLAIN HOW DID YOU GET YOUR ANSWER?

OpenStudy (anonymous):

http://www.analyzemath.com/line/line.htm

OpenStudy (anonymous):

x and y Intercepts of the Graph of a Line We now explore more general equations of lines with equations ax + by = c, where a is not equal to zero and b is not equal to zero. The x intercept is found by setting y = 0 in the above equation and solve for x. ax + b(0) = c x = c/a Hence, the x intercept is at (c/a , 0). The y intercept is found by setting x = 0 in the above equation and solve for x. a(0) + by = c y = c/b Hence, the y intercept is at (0 , c/b).

OpenStudy (anonymous):

did you understood it ?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

then it is ok

OpenStudy (anonymous):

wait, actually I do now, thanks ! :)